OpenStudy (df001):

For the value of Θ, use half angle formula find Sin(1/2)Θ Θ=300 degrees

OpenStudy (anonymous):

cosA=1- 2sin(A/2)^2 now A=300 then cosA=1/2 therfore 1/2=1-2sin(A/2)^2 =>2sin(A/2)^2=1-1/2 =>sin(A/2)^2=1/4 =>sin(A/2)=1/2 as A=150 and it is in the second quadrant